Accelerator Update for August 20, 2008 - August 22, 2008
Wednesday August 20
The day shift began with Operations monitoring the Tevatron (TeV) store 6368 (96.12E30) and the Recycler (RR) stash (247.5E10), with the Antiproton Source (Pbar) stacking (36.6E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam.
At 8:50 AM, an expert reported that the TeV sector A0 kicker room air conditioner probably had a broken belt. With permission, they accessed the top of the TeV berm and replaced the belt. The room temperature returned to normal.
A Pbar expert reported at 9:49 AM that he was raising the Lithium Lens gradient voltage to its normal operating value of 2600 volts.
A MI expert began a study at 10:59 AM.
Operations terminated store 6368 at 11:59 AM.
Operations prepared D0 and CDF for access at 12:20 PM.
Operations prepared Pbar for access at 12:28 PM.
Operations prepared the TeV for access at 12:46 PM.
Operations prepared the TeV F sector for access at 12:59 PM.
Operations prepared the NuMI absorber area and alcoves for access at 1:48 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons three times during the day shift.
D0 came out of access at 4:28 PM.
Pbar came out of access at 5:16 PM.
Operations established beam to MiniBooNE at 5:22 PM.
Operations established beam to NuMI at 5:37 PM.
At 5:52 PM, Operations reported a problem with stacking. Operators investigated and found that a Pbar power supply (V701) would not turn on. Experts were called in to investigate.
The TeV was out of access by 5:53 PM.
At 7:41 PM, the NuMI beam permit tripped off due to a current sag on a quadrupole power supply (E:QF107). Everything reset.
Operations began the TeV turn on at 9:31 PM.
At 9:46 PM, E:QF107 tripped off and caused the NuMI beam permit to also trip. Everything reset.
Power supply experts reported at 10:32 PM that they had completed their work on the Pbar power supply (V701) and brought it back on line.
Operations established beam to the Accumulator at 10:40 PM, which then caused the Lithium Lens to trip off. Operators investigated, contacted a Pbar expert, and then called in power supply experts.
Operations put the TeV through a wet squeeze at 11:33 PM.
Operations put the TeV into shot setup at 11:54 PM
Thursday August 21
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (342.3E10), with the Pbar not stacking due to a lens problem, and with the MiniBooNE and NuMI experiments taking beam.
Operations established store 6371 at 1:26 AM with an initial luminosity of 272.21E30.
A Pbar expert reported that the Lithium Lens power supply was fixed. Operations established beam to Pbar at 5:57 AM.
A water systems expert was asked to check the LCW hoses for Linac Klystron station #7. He reported at 8:48 AM that the hoses for this Klystron station as well as the others are old and need replacing.
At 11:22 AM, the P2 transfer line beam permit tripped off due to QF12. Everything reset.
A MI expert began a study period at 1:51 PM.
At 2:46 PM, a quadrupole magnet power supply (QF107) tripped off for the first of four times in 45 minutes and each trip caused the NuMI beam permit to also trip. Everything reset.
FESS reported at 3:18 PM that they had repaired the ICW leak in the NS3 parking lot.
Operations reported at 4:02 PM that the Recycler had stashed six times during the day shift.
At 4:13 PM, Operations reported a MIRF Watchdog trip, which caused all the MIRF stations to trip off. Everything reset.
At 4:28 PM, Operations reported another MIRF Watchdog trip. Operations contacted an expert.
The MI study ended at 4:36 PM.
An expert reported at 4:47 PM that he found a timing jitter on QF107 and QF12. He changed their controls card.
Operations disabled the MTest beam at 6:09 PM.
At 11:07 PM, KRF6 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Friday August 22
The midnight shift began with Operations monitoring store 6371 and the stash (349.2E10), with Pbar stacking, and with the MiniBooNE and NuMI experiments taking beam.
Operations terminated store 6371 at 12:30 AM.
Operations put the TeV through a wet squeeze at 1:07 AM.
Operations put the TeV into shot setup at 1:51 AM.
At 2:06 AM, BRF11 tripped off. It reset.
Operations established store 6373 at 3:42 AM with an initial luminosity of 264.07E30.
Operations disabled the MTest beam around 5:55 AM.
At 5:56 AM, LRF4 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 08/22/2008 email Fermilab|
|Security, Privacy, Legal|